# On nonlinear transitions, minimal seeds and exact solutions for the geodynamo

### v1

@@ 4 Conclusions
Motivated by the search for strong branch solutions of the geodynamo, we have developed the code to perform adjoint-based optimal control on a convection-driven, magnetohydrodynamic flow in a rotating spherical shell, using the Dedalus environment. We have revisited the well-studied dynamo benchmark of Christensen et al. (2001) as an example case to systematically identify minimal magnetic perturbations (in the sense of least magnetic energy) that nonlinearly trigger a convective dynamo, relying on transient growth optimisation.
@@
Whereas in the systems considered by Mannix et al. (2022), maximising the time-integrated magnetic energy over a fraction of the ohmic timescale was found suitable to robustly identify dynamo solutions and their minimal seeds, in the present system this approach appears less reliable. Indeed, the optimisation procedure identified multiple magnetic conditions that (locally) maximise magnetic energy growth but do not trigger a sustained dynamo at longer times. We are facing here two practical difficulties: firstly, the optimisation problem is highly nonconvex; secondly, we are restricted to relatively short time horizons for the optimisation. Indeed we found optimisation to become highly sensitive to small changes in the initial conditions as soon as the target time approached $T_{\textrm{opt}}=0.4$, rendering convergence difficult; this observation is however readily explained by the fact that the system was approaching the (unstable) edge state by this time. Moreover, long-time optimisations yield less accurate gradients — especially since for convenience we use a differentiate-then-discretise approach to integrate the adjoint equations (Mannix et al., 2024), so that increasing $T_{\textrm{opt}}$ past 0.4 is not realistic. Yet, the optimisation question we are truly interested in is best investigated over (very) long time horizons: indeed the very definition of dynamo action is the existence of an initial condition such that magnetic energy does not vanish as time goes to infinity! Were $T_{\textrm{opt}}$ allowed to become long enough while maximising the magnetic energy, then all initial conditions yielding failed dynamos would essentially yield negligible objective functions, while any local maximum would have to be a dynamo seed. Whereas for short target times, maximising the total magnetic energy does not penalise local extrema corresponding to initial conditions that trigger strong transient growth ultimately followed by decay. In other words, we need for practical reasons to devise a cost functional that, when extremised over relatively short time horizon, acts as a robust proxy for discriminating long times evolution.
@@
This situation is reminiscent of a similar issue encountered in optimal mixing problems, where multiscale cost functions can act as shorter-target time proxies for long-time flow evolution. Specifically, minimising Sobolev norms of negative indexes called the mix-norms (Mathew et al., 2005; Doering and Thiffeault, 2006) by a short target time has been shown to effectively yield the same results (in terms of both the optimal perturbations and the achieved objective functions) as minimising the flow variance over long target times, whether in unstratified (Foures et al., 2014; Heffernan and Caulfield, 2022) or stratified fluids (Marcotte and Caulfield, 2018). Mix-norms thus provide a computationally efficient way of identifying the long-term, physical mixing processes while working with affordable target times.
@@
Moreover, the identification of travelling waves solutions using a Newton-hookstep algorithm (Willis, 2019b) confirms that the evolution initiated by these optimal dynamo seeds brings the system close to an unstable magnetohydrodynamic solution within the time horizon we used for the computation of the minimal seed. From that unstable state, the fate of the system is determined by the energy budget allowed for the initial seeds, travelling away toward either the dynamo attractor (which also corresponds to the linearly stable, magnetised travelling wave solution identified by the Newton-hookstep algorithm) or a non-dynamo one (the linearly stable, purely hydrodynamic travelling wave solution).
@@
Therefore, we can now form a picture of the dynamical landscape of the various basins of attraction at our working parameters, which summarises the findings of the present paper. Figure 7 shows the trajectory initiated by different initial conditions in the kinetic-magnetic energy phase space, along with the travelling waves solutions (hydrodynamic TW0, stable MHD TW1, unstable MHD TW2). Whereas the Christensen et al. (2001) benchmark initial condition (rescaled here to the largest energy at which it can no longer trigger a dynamo) approaches the unstable edge state TW1 in a (somewhat) inefficient way, undergoing repeated energy transfers between the kinetic, magnetic and thermal energy reservoirs, the less energetic, minimal dynamo seed yields an important amplification of the magnetic energy by maintaining a stronger convection at early times, while energy transfers at later times are largely suppressed. This strategy rapidly propels the system to the vicinity of the edge state, from which it is repelled away towards the dynamo attractor or (when the initial energy budget is too small) the purely hydrodynamic one.
@@
We conclude by returning to our motivation. What is the relevance of our approach for studies of the geodynamo? Finding minimal seeds and exact nonlinear solutions will be important in promoting our understanding in a field where it is impossible to perform numerical calculations in the correct parameter regime. We intend to build on this work by investigating how the minimal seeds and exact nonlinear solutions change as the underlying parameters are varied. For example, how does the minimal energy of the magnetic field perturbation change as one varies the parameters along the dynamo path (Aubert et al., 2017) taking one to the distinguished limit of Dormy (2016). Does the energy required to take one to a fully nonlinear strong field solution decrease as the Ekman number is decreased; this behaviour would be reminiscent of the behaviour with Reynolds number of the energy of the minimal seed in wall-bounded shear flows (see e.g. Parente et al., 2022). If this were the case then even a tiny magnetic energy perturbation would be enough to recover the dynamo if it entered a non-dynamo state — even if the hydrodynamic solution were formally linearly stable. Furthermore, tracing the behaviour of exact solutions as parameters are varied is a very efficient way of examining whether (admittedly specialised) solutions could maintain balance as parameters are varied. The bifurcation structure of these solutions (particularly the presence or absence of global bifurcations) may also give hints about mechanisms for achieving reversal of solutions whilst maintaining the correct balance.
@@ Acknowledgements
This work was undertaken on ARC4, part of the High Performance Computing facilities at the University of Leeds, UK. CSS would like to acknowledge Dr Jacob Page for helpful discussions on the Newton-hookstep algorithm.
@@ A.1 Adjoints of vector calculus operations
To adjoint the scalar Laplacian we can use our results for the divergence and gradient to obtain
@@
The vector Laplacian can also be adjointed from the rules we have derived so far, by writing $\nabla^{2}\bm{y}=\nabla(\nabla\cdot\bm{y})-\nabla\times(\nabla\times\bm{y})$, yielding
@@ A.2 The adjoint equations
With the adjoints of general vector calculus operations now derived, we can proceed to adjoint our equations by cancelling all first variations of the Lagrangian (11). Taking the first variation with respect to the adjoint state ensures that the governing equations are satisfied. The first variation with respect to the state $\bm{q}$, gives
$\nabla_{\bm{q}}\mathcal{L}\cdot\delta\bm{q}=\left[\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}},\delta\bm{q}\right]+\left[\mathbf{M}^{T}\frac{\partial\bm{q}^{\dagger}}{\partial t},\delta\bm{q}\right]+\left[\bm{q}^{\dagger},\frac{\partial\bm{\mathcal{N}}(\bm{q})}{\partial\bm{q}}\delta\bm{q}\right],$ (41)
@@
$\nabla_{\bm{q}}\mathcal{L}\cdot\delta\bm{q}=\left[\mathbf{M}^{T}\frac{\partial\bm{q}^{\dagger}}{\partial t}+\bm{\mathcal{A}}^{\dagger}\bm{q}^{\dagger}+\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}},\delta\bm{q}\right].$ (43)
@@
$-\mathbf{M}^{T}\frac{\partial\bm{q}^{\dagger}}{\partial t}=\bm{\mathcal{A}}^{\dagger}\bm{q}^{\dagger}+\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}}.$ (44)
In order to find the adjoint operator $\bm{\mathcal{A}}^{\dagger}$, equation (42) shows us that we must first linearise the governing equations. Integration-by-parts using the rules in table 2 then yields the adjoint operator. For our equations, this is done using the steps outlined by Chen et al. [2018], where the governing equations are rewritten in a form in which Ohm’s law (where Ampère’s law has been used to relate the magnetic field to the current)
@@
appear separately. In other words, equations for the electric field $\bm{E}$ and magnetic field are constrained separately using adjoint variables $\bm{E}^{\dagger}$ and $\bm{b}^{\dagger}$, respectively. As in Chen et al. [2018], $\sigma$ is a relative electrical conductivity which takes the value $1$ for $\bm{x}\in\mathcal{V}_{s}$ and 00 for $\bm{x}\in\mathbb{R}^{3}\backslash\mathcal{V}_{s}$. This allows the adjoint formulation to more easily incorporate electrically insulating inner and outer regions, where the insulating boundary condition is simply continuity of the radial part of the magnetic field and tangential part of the electric field at the boundaries between $\mathcal{V}_{s}$ and the insulating regions.
@@
These equations match the form of the equations contained in [Mannix et al., 2022], and also [Chen et al., 2018, Holdenried-Chernoff et al., 2019] if only the induction equation with a specified velocity field is considered, but with extra terms stemming from the temperature variable. We see that the adjoint equations currently have more unknowns than equations due to the inclusion of the $\nabla\Pi^{\dagger}$ term in the adjoint magnetic field equation. As in the study of Chen et al. [2018], we see that the adjoint equations imply a gauge freedom for the adjoint magnetic field $\bm{b}^{\dagger}$. As it is only the curl of the adjoint magnetic field that occurs, setting $\bm{b}^{\dagger}\rightarrow\bm{b}^{\dagger}+\nabla\chi$ does not change the equations. It turns out to be convenient to then fix this gauge freedom by choosing $\bm{b}^{\dagger}$ to be solonoidal, i.e. we solve (14) together with $\nabla\cdot\bm{b}^{\dagger}=0$, which determines the value of $\Pi^{\dagger}$. This gauge choice has two main benefits; firstly that the update condition (17) is now guaranteed to give an updated magnetic field that is divergence free, and secondly the boundary conditions for the adjoint magnetic field will be the same as for the direct magnetic field (see Chen et al. [2018], for example). Cancelling the boundary terms that arise from integration-by-parts of terms involving the velocity and temperature fields gives the adjoint boundary conditions displayed in equation (16).
@@ Appendix B Cost functionals
For each cost functional considered we need to be able to find its contribution to the adjoint equation. For optimising the total magnetic energy
$\mathcal{J}=\frac{1}{2\textit{Ek}\textit{Pm}}\iiint_{\mathcal{V}_{s}}\bm{B}\cdot\bm{B}\;\textrm{d}V,$ (49)
giving
@@
Similarly, for optimising the axisymmetric energy
$\mathcal{J}=\frac{1}{2\textit{Ek}\textit{Pm}}\iiint_{\mathcal{V}_{s}}\bar{\bm{B}}\cdot\bar{\bm{B}}\;\textrm{d}V,$ (51)
we obtain
@@ Appendix C Convergence
where $\bm{g}_{\perp}$ is the Riemannian gradient that does not change the norm of $\bm{B}_{0}$. The cost functional rapidly increases at first, as the large scale structure of the initial magnetic field undergoes rapid changes from its random initial guess. After approximately 20 iterations the cost functional tends toward a constant value as minor, mostly small scale adjustments occur, while the relative residual steadily decreases.
(a) Convergence of the optimisation procedure with the total energy cost functional.
(b) Convergence of the optimisation procedure with the $m=0$-based cost functional.

### v2

@@ 4 Conclusions
Motivated by the search for strong branch solutions of the geodynamo, we have developed the code to perform adjoint-based optimal control on a convection-driven, magnetohydrodynamic flow in a rotating spherical shell, using the Dedalus environment. We have revisited the well-studied dynamo benchmark of Christensen et al., (2001) as an example case to systematically identify minimal magnetic perturbations (in the sense of least magnetic energy) that nonlinearly trigger a convective dynamo, relying on transient growth optimisation.
@@
Whereas in the systems considered by Mannix et al., (2022), maximising the time-integrated magnetic energy over a fraction of the ohmic timescale was found suitable to robustly identify dynamo solutions and their minimal seeds, in the present system this approach appears less reliable. Indeed, the optimisation procedure identified multiple magnetic conditions that (locally) maximise magnetic energy growth but do not trigger a sustained dynamo at longer times. We are facing here two practical difficulties: firstly, the optimisation problem is highly nonconvex; secondly, we are restricted to relatively short time horizons for the optimisation. Indeed we found optimisation to become highly sensitive to small changes in the initial conditions as the target time approached $t_{\textrm{opt}}=0.4$, rendering convergence difficult; this observation is however readily explained by the fact that the system was approaching the (unstable) edge state by this time. Moreover, our use of a differentiate-then-discretise approach to solve the adjoint equations (Mannix et al.,, 2024) causes our gradient estimate to deteriorate as the time horizon is increased. Yet, the optimisation question we are truly interested in is best investigated over (very) long time horizons: indeed the very definition of dynamo action is the existence of an initial condition such that magnetic energy does not vanish as time goes to infinity! Were $t_{\textrm{opt}}$ allowed to become long enough while maximising the magnetic energy, then all initial conditions yielding failed dynamos would essentially yield negligible objective functions, while any local maximum would have to be a dynamo seed. Whereas for short target times, maximising the total magnetic energy does not penalise local extrema corresponding to initial conditions that trigger strong transient growth ultimately followed by decay. In other words, we need for practical reasons to devise a cost functional that, when extremised over relatively short time horizon, acts as a robust proxy for discriminating long times evolution.
@@
This situation is reminiscent of a similar issue encountered in optimal mixing problems, where multiscale cost functions can act as shorter-target time proxies for long-time flow evolution. Specifically, minimising Sobolev norms of negative indexes called the mix-norms (Mathew et al.,, 2005; Doering and Thiffeault,, 2006) by a short target time has been shown to effectively yield the same results (in terms of both the optimal perturbations and the achieved objective functions) as minimising the flow variance over long target times, whether in unstratified (Foures et al.,, 2014; Heffernan and Caulfield,, 2022) or stratified fluids (Marcotte and Caulfield,, 2018). Mix-norms thus provide a computationally efficient way of identifying the long-term, physical mixing processes while working with affordable target times.
@@
Moreover, the identification of travelling waves solutions using a Newton-hookstep algorithm (Willis, 2019a, ) confirms that the evolution initiated by these optimal dynamo seeds brings the system close to an unstable magnetohydrodynamic solution within the time horizon we used for the computation of the minimal seed. From that unstable state, the fate of the system is determined by the energy budget allowed for the initial seeds, travelling away toward either the dynamo attractor (which also corresponds to the linearly stable, magnetised travelling wave solution identified by the Newton-hookstep algorithm) or a non-dynamo one (the linearly stable, purely hydrodynamic travelling wave solution).
@@
Therefore, we can now form a picture of the dynamical landscape of the various basins of attraction at our working parameters, which summarises the findings of the present paper. Figure 7 shows the trajectory initiated by different initial conditions in the kinetic-magnetic energy phase space, along with the travelling waves solutions (hydrodynamic TW0, stable MHD TW1, unstable MHD TW2). Whereas the Christensen et al., (2001) benchmark initial condition (rescaled here to the largest energy at which it can no longer trigger a dynamo) approaches the unstable edge state TW1 in a (somewhat) inefficient way, undergoing repeated energy transfers between the kinetic, magnetic and thermal energy reservoirs, the less energetic, minimal dynamo seed yields an important amplification of the magnetic energy by maintaining a stronger convection at early times, while energy transfers at later times are largely suppressed. This strategy rapidly propels the system to the vicinity of the edge state, from which it is repelled away towards the dynamo attractor or (when the initial energy budget is too small) the purely hydrodynamic one.
@@
We conclude by returning to our motivation. What is the relevance of our approach for studies of the geodynamo? Finding minimal seeds and exact nonlinear solutions will be important in promoting our understanding in a field where it is impossible to perform numerical calculations in the correct parameter regime. We intend to build on this work by investigating how the minimal seeds and exact nonlinear solutions change as the underlying parameters are varied. For example, how does the minimal energy of the magnetic field perturbation change as one varies the parameters along the dynamo path (Aubert et al.,, 2017) taking one to the distinguished limit of Dormy, (2016). Does the energy required to take one to a fully nonlinear strong field solution decrease as the Ekman number is decreased; this behaviour would be reminiscent of the behaviour with Reynolds number of the energy of the minimal seed in wall-bounded shear flows (see e.g. Duguet et al.,, 2013). If this were the case then even a tiny magnetic energy perturbation would be enough to recover the dynamo if it entered a non-dynamo state — even if the hydrodynamic solution were formally linearly stable. Furthermore, tracing the behaviour of exact solutions as parameters are varied is a very efficient way of examining whether (admittedly specialised) solutions could maintain balance as parameters are varied. The bifurcation structure of these solutions (particularly the presence or absence of global bifurcations) may also give hints about mechanisms for achieving reversal of solutions whilst maintaining the correct balance.
@@ Acknowledgements
CSS would like to acknowledge Jacob Page for helpful discussions on the Newton-hookstep algorithm. The authors thank Yohann Duguet for providing valuable feedback on the manuscript. This work was undertaken on ARC4, part of the High Performance Computing facilities at the University of Leeds, UK.
@@ A.1 Adjoints of vector calculus operations
To find the adjoint of the scalar Laplacian we can use our results for the divergence and gradient to obtain
@@
The adjoint of the vector Laplacian can also be found from the rules we have derived so far, by writing $\nabla^{2}\bm{y}=\nabla(\nabla\cdot\bm{y})-\nabla\times(\nabla\times\bm{y})$, yielding
@@ A.2 The adjoint equations
With the adjoints of general vector calculus operations now derived, we can proceed to find the adjoint our equations by cancelling all first variations of the Lagrangian (10). Taking the first variation with respect to the adjoint state ensures that the governing equations are satisfied. The first variation with respect to the state $\bm{q}$, gives
$\nabla_{\bm{q}}\mathcal{L}\cdot\delta\bm{q}=\left[\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}},\delta\bm{q}\right]+\left[\mathbf{M}^{\mathrm{T}}\frac{\partial\bm{q}^{\dagger}}{\partial t},\delta\bm{q}\right]+\left[\bm{q}^{\dagger},\frac{\partial\bm{\mathcal{N}}(\bm{q})}{\partial\bm{q}}\delta\bm{q}\right],$ (38)
@@
$\nabla_{\bm{q}}\mathcal{L}\cdot\delta\bm{q}=\left[\mathbf{M}^{\mathrm{T}}\frac{\partial\bm{q}^{\dagger}}{\partial t}+\bm{\mathcal{A}}^{\dagger}\bm{q}^{\dagger}+\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}},\delta\bm{q}\right].$ (40)
@@
$-\mathbf{M}^{\mathrm{T}}\frac{\partial\bm{q}^{\dagger}}{\partial t}=\bm{\mathcal{A}}^{\dagger}\bm{q}^{\dagger}+\frac{\partial\mathcal{J}_{I}}{\partial\bm{q}}.$ (41)
In order to find the adjoint operator $\bm{\mathcal{A}}^{\dagger}$, equation (39) shows us that we must first linearise the governing equations. Integration-by-parts using the rules in table 2 then yields the adjoint operator. For our equations, this is done using the steps outlined by Chen et al., (2018), where the governing equations are rewritten in a form in which Ohm’s law (where Ampère’s law has been used to relate the magnetic field to the current)
@@
appear separately. In other words, equations for the electric field $\bm{E}$ and magnetic field are constrained separately using adjoint variables $\bm{E}^{\dagger}$ and $\bm{b}^{\dagger}$, respectively. As in Chen et al., (2018), $\sigma$ is a relative electrical conductivity which takes the value $1$ for $\bm{x}\in\mathcal{V}_{s}$ and 00 for $\bm{x}\in\mathbb{R}^{3}\backslash\mathcal{V}_{s}$. This allows the adjoint formulation to more easily incorporate electrically insulating inner and outer regions, where the insulating boundary condition is simply continuity of the radial part of the magnetic field and tangential part of the electric field at the boundaries between $\mathcal{V}_{s}$ and the insulating regions. Additionally, we introduce a Lagrange multiplier $\Pi^{\dagger}$, which constraints the solenoidal condition $\nabla\cdot\bm{B}=0$.
@@
These equations match the form of the equations contained in (Mannix et al.,, 2022), and also (Chen et al.,, 2018; Holdenried-Chernoff et al.,, 2019) if only the induction equation with a specified velocity field is considered, but with extra terms stemming from the temperature variable. We see that the adjoint equations currently have more unknowns than equations due to the inclusion of the $\nabla\Pi^{\dagger}$ term in the adjoint magnetic field equation. As in the study of Chen et al., (2018), we see that the adjoint equations imply a gauge freedom for the adjoint magnetic field $\bm{b}^{\dagger}$. As it is only the curl of the adjoint magnetic field that occurs, setting $\bm{b}^{\dagger}\rightarrow\bm{b}^{\dagger}+\nabla\chi$ does not change the equations. This gauge freedom is a consequence of the direct equations being overdetermined for $\bm{B}$, with the induction equation naturally preserving the divergence of $\bm{B}$ which is physically always initialised with a divergence free field. It turns out to be convenient to then fix this gauge freedom by choosing $\bm{b}^{\dagger}$ to be solonoidal, i.e. we solve (14) together with $\nabla\cdot\bm{b}^{\dagger}=0$, which determines the value of $\Pi^{\dagger}$. This gauge choice has two main benefits; firstly that the update condition (13) is now guaranteed to give an updated magnetic field that is divergence free, and secondly the boundary conditions for the adjoint magnetic field will be the same as for the direct magnetic field (see Chen et al., (2018), for example). Cancelling the boundary terms that arise from integration-by-parts of terms involving the velocity and temperature fields gives the adjoint boundary conditions.
@@ Appendix B Cost functionals
For each cost functional considered we need to be able to find its contribution to the adjoint equation. For optimising the total magnetic energy (30), we obtain
@@
Similarly, for optimising the axisymmetric energy (31) we find
@@ Appendix C Convergence
where $\bm{g}_{\perp}$ is the Riemannian gradient that does not change the norm of $\bm{B}_{0}$. The cost functional rapidly increases at first, as the large scale structure of the initial magnetic field undergoes rapid changes from its random initial guess. We perform the optimisation procedure until either the residual is less than $10^{-3}$ or 120 iterations have occurred. After approximately 20 iterations the cost functional tends toward a constant value as minor, mostly small-scale adjustments occur, while the relative residual steadily decreases.
(a) Total energy cost functional.
(b) Axisymmetric energy cost functional.